The spacecraft travels at a speed of 0.737c relative to Earth. A motorcycle race is taking place on the surface of the Earth. The passengers in the ship measure the duration of the race that takes place on the surface of the Earth and that time on their watch is 15.25 s. What is the actual duration of the race?


1
Expert's answer
2021-03-04T15:19:25-0500
"t'=\frac{t}{\sqrt{1-\frac{v^2}{c^2}}}\\
t'=\frac{15.25}{\sqrt{1-0.737^2}}=22.56\ s"